# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v2

@@ 9.2 Evaluation Metrics
The evaluation metrics for missing data imputation can be organized into four groups, shown in Table 7. The first group is composed of various imputation errors, based on the difference between the imputed value and the true value. Let $x_{i,j}$ and $\hat{x}_{i,j}$ be the true value and imputed value respectively, where $(i,j)\in\Omega$ and $\Omega$ denotes the index set of all missing values. The mean absolute error is defined as $\text{MAE}:=\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}|x_{i,j}-\hat{x}_{i,j}|$. Similarly, the root mean square error is $\text{RMSE}:=\sqrt{\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}(x_{i,j}-\hat{x}_{i,j})^{2}}$. MAE is less sensitive to outliers than RMSE. It is worth noting that both MAE and RMSE cannot be compared across different datasets unless the data in each dataset are standardized. For instance, when the data are standardized into $[0,1]$, we can compare the imputation difficulty between two datasets. One common limitation of MAE and RMSE is that they cannot indicate how large the imputation error is relative to the ground truth; although standardization to $[0,1]$ helps, it is still sensitive to outliers. Therefore, we should use the relative MAE and relative RMSE, defined as $r\text{MAE}:=\sum_{(i,j)\in\Omega}|x_{i,j}-\hat{x}_{i,j}|/\sum_{(i,j)\in\Omega}|x_{i,j}|$ and $r\text{RMSE}:=\sqrt{\sum_{(i,j)\in\Omega}(x_{i,j}-\hat{x}_{i,j})^{2}/\sum_{(i,j)\in\Omega}x_{i,j}^{2}}$, respectively.
The second group is based on distance metrics between distributions such as the Wasserstein distance (WD) (villani2008optimal) and maximum mean discrepancy (MMD) (gretton2012kernel). Let $\mathbf{X}\in\mathbb{R}^{d\times n_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{d\times n_{y}}$ be two discrete distributions in $\mathbb{R}^{d}$, WD2 (Wasserstein-2 distance) is calculated by
@@
As can be seen, the computations of WD and MMD are much more time-consuming than the imputation errors, but they can be applied to cases where the ground truth of the missing values is not available. For example, let $\tilde{\mathbf{X}}_{a}$ be an incomplete data matrix, for which the ground truth of the missing values is not available. Let $\hat{\mathbf{X}}_{a}$ be the imputed data matrix given by some imputation method. If there is another complete data matrix ${\mathbf{X}}_{b}$ from the same distribution as $\mathbf{X}_{a}$, we can measure the WD between $\hat{\mathbf{X}}_{a}$ and $\mathbf{X}_{b}$, where a smaller WD means a better imputation performance. WD has been utilized in a few papers such as (muzellec2020missing; jarrett2022hyperimpute).
The third group of evaluation metrics assesses performance on downstream tasks—specifically regression, classification, and clustering. For regression, the most common measures are the Mean Absolute Error MAE and RMSE. For classification, standard metrics include accuracy, precision, recall, F1-score, AUROC, and AUPRC; grandini2020metrics provided a comprehensive survey of these measures. Clustering metrics fall into two families: internal indices, which do not require ground-truth labels (e.g., the Davies–Bouldin index and Silhouette coefficient), and external indices, which do (e.g., clustering accuracy, normalized mutual information (NMI), and the adjusted rand index (ARI)).
@@ 10.1 Model Selection and Hyperparameter Tuning for Missing Data Imputation
Method or model selection for missing data imputation is challenging. Almost every imputation method has at least one hyperparameter and the best values for the hyperparameters vary across different datasets. For each dataset, we need to split it into two subsets, one for training and the other for validation, to find the best hyperparameters. This will be time-consuming when the number of hyperparameters is large and the imputation method has a high-time complexity. For instance, for a deep learning-based imputation method, we need to determine the number of layers of the neural network, the width of each layer, the learning rate, the number of iterations, the coefficient of weight decay, and algorithm-specific hyperparameters; if each has 4 candidate values, we need to use grid search over at least 4096 combinations, meaning training 4096 neural networks, which is extremely costly when the dataset is large. To accelerate the hyperparameter tuning, we may use more effective strategies such as Bayesian optimization (snoek2012practical) or other automated machine learning (AutoML) techniques (hutter2019automated) instead of grid search. Although some works, such as (jarrett2022hyperimpute), have used AutoML to tune the hyperparameters of imputation methods, the related studies are limited, and there is a large room for improvement. Future work may consider the following points:
@@ 10.2 Privacy Protection in Missing Data Imputation
Data privacy is the cornerstone of trustworthy data science: without it, individuals can be re-identified from supposedly anonymized models, exposing medical records, financial details, or location histories. For missing data imputation, we also need to consider the problems of data leakage and privacy protection. Privacy-preserving techniques like differential privacy (dwork2006differential) and federated learning (kairouz2021advances) enable organizations to unlock collective insights across silos—think hospitals training shared cancer-detection models without ever moving patient data—turning privacy from a legal checkbox into a competitive advantage. In federated-learning scenarios, the training data are distributed across multiple clients, and any sharing or transmission between clients—or between a client and the central server, if one exists—is forbidden, thereby protecting data privacy to some extent. This protection is often further strengthened by differential privacy, with a trade-off between accuracy and privacy.
Recently, there have been a few studies considering federated learning for missing data imputation (10.14778/3503585.3503598; abbasi2023fast; lian2025federated; fang2025msfgan; 10.1016/j.eswa.2025.126543), but there are still a few challenges.
  • It is difficult to extend nonparametric imputation methods such as kNN and tree-based methods to federated learning scenarios (zhang2024fedknn; Li_Wen_He_2020). For example, regarding kNN, we cannot directly compute the distance between two data points located on different clients.
@@ 10.4 Pretrained and Universal Models for Missing Data Imputation
In recent years, pretrained models, such as BERT (devlin2019bert), CLIP (radford2021learning), and GPTs (achiam2023gpt), have become the centerpiece of AI research, delivering breakthrough performance on a wide spectrum of real-world problems. A pretrained model is a large neural network whose parameters have already been learned from massive, general-domain data—text, images, audio, protein sequences, or code—so it arrives packed with broad patterns and world knowledge; instead of training from scratch, practitioners simply fine-tune or prompt it with a handful of examples, slashing data, compute, and time by orders of magnitude. It is natural to consider using pretrained models in missing data imputation or even establishing a universal model for missing data imputation. Here, a universal model for missing data imputation refers to one that, once well trained, can impute missing values for any dataset—or at least a broad class of datasets from different domains—without requiring further tuning or with only light fine-tuning.
As discussed in Section 5.7, a few scholars have explored the use of LLMs to assist in missing data imputation (huang2024missing; hayat2024context; fang2025spatiotemporal) and achieved some inspiring results in general missing data imputation tasks. However, there are still limitations:
@@ 11 Conclusion
Acknowledgements

### v3

@@ 9.2 Evaluation Metrics
The evaluation metrics for missing data imputation can be organized into four groups, shown in Table 7. The first group is composed of various imputation errors, based on the difference between the imputed value and the true value. Let $x_{i,j}$ and $\hat{x}_{i,j}$ be the true value and imputed value respectively, where $(i,j)\in\Omega$ and $\Omega$ denotes the index set of all missing values. The mean absolute error is defined as $\text{MAE}:=\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}|x_{i,j}-\hat{x}_{i,j}|$. Similarly, the root mean square error is $\text{RMSE}:=\sqrt{\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}(x_{i,j}-\hat{x}_{i,j})^{2}}$. MAE is less sensitive to outliers than RMSE. It is worth noting that both MAE and RMSE cannot be compared across different datasets unless the data in each dataset are standardized. For instance, when the data are standardized into $[0,1]$, we can compare the imputation difficulty between two datasets. One common limitation of MAE and RMSE is that they cannot indicate how large the imputation error is relative to the ground truth; although standardization to $[0,1]$ helps, it is still sensitive to outliers. Therefore, we should use the relative MAE and relative RMSE, defined as $r\text{MAE}:=\sum_{(i,j)\in\Omega}|x_{i,j}-\hat{x}_{i,j}|/\sum_{(i,j)\in\Omega}|x_{i,j}|$ and $r\text{RMSE}:=\sqrt{\sum_{(i,j)\in\Omega}(x_{i,j}-\hat{x}_{i,j})^{2}/\sum_{(i,j)\in\Omega}x_{i,j}^{2}}$, respectively. A practical consideration for evaluating imputation methods on real-world tabular data is the handling of heterogeneous feature types. While metrics like RMSE and MAE are standard for continuous variables, categorical, ordinal, and binary variables necessitate different evaluation criteria (e.g., accuracy, F1-score, or weighted Cohen’s kappa). Therefore, a complete benchmarking framework for such data should employ a hybrid evaluation strategy that combines type-specific metrics or their aggregated scores to reflect overall imputation quality across all feature types.
The second group is based on distance metrics between distributions such as the Wasserstein distance (WD) (Villani and others, 2008) and maximum mean discrepancy (MMD) (Gretton et al., 2012). Let $\mathbf{X}\in\mathbb{R}^{d\times n_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{d\times n_{y}}$ be two discrete distributions in $\mathbb{R}^{d}$, WD2 (Wasserstein-2 distance) is calculated by
@@
As can be seen, the computations of WD and MMD are much more time-consuming than the imputation errors, but they can be applied to cases where the ground truth of the missing values is not available. For example, let $\tilde{\mathbf{X}}_{a}$ be an incomplete data matrix, for which the ground truth of the missing values is not available. Let $\hat{\mathbf{X}}_{a}$ be the imputed data matrix given by some imputation method. If there is another complete data matrix ${\mathbf{X}}_{b}$ from the same distribution as $\mathbf{X}_{a}$, we can measure the WD between $\hat{\mathbf{X}}_{a}$ and $\mathbf{X}_{b}$, where a smaller WD means a better imputation performance. WD has been utilized in a few papers such as (Muzellec et al., 2020; Jarrett et al., 2022).
The third group of evaluation metrics assesses performance on downstream tasks—specifically regression, classification, and clustering. For regression, the most common measures are the Mean Absolute Error MAE and RMSE. For classification, standard metrics include accuracy, precision, recall, F1-score, AUROC, and AUPRC; Grandini et al. (2020) provided a comprehensive survey of these measures. Clustering metrics fall into two families: internal indices, which do not require ground-truth labels (e.g., the Davies–Bouldin index and Silhouette coefficient), and external indices, which do (e.g., clustering accuracy, normalized mutual information (NMI), and the adjusted rand index (ARI)).
@@ 10.1 Model Selection and Hyperparameter Tuning for Missing Data Imputation
Method or model selection for missing data imputation is challenging. Almost every imputation method has at least one hyperparameter and the best values for the hyperparameters vary across different datasets. For each dataset, we need to split it into two subsets, one for training and the other for validation, to find the best hyperparameters. This will be time-consuming when the number of hyperparameters is large and the imputation method has a high-time complexity. For instance, for a deep learning-based imputation method, we need to determine the number of layers of the neural network, the width of each layer, the learning rate, the number of iterations, the coefficient of weight decay, and algorithm-specific hyperparameters; if each has 4 candidate values, we need to use grid search over at least 4096 combinations, meaning training 4096 neural networks, which is extremely costly when the dataset is large. To accelerate the hyperparameter tuning, we may use more effective strategies such as Bayesian optimization (Snoek et al., 2012) or other automated machine learning (AutoML) techniques (Hutter et al., 2019) instead of grid search. Although some works, such as (Jarrett et al., 2022), have used AutoML to tune the hyperparameters of imputation methods, the related studies are limited, and there is a large room for improvement. Future work may consider the following points:
@@ 10.2 Privacy Protection in Missing Data Imputation
Data privacy is the cornerstone of trustworthy data science: without it, individuals can be re-identified from supposedly anonymized models, exposing medical records, financial details, or location histories. For missing data imputation, we also need to consider the problems of data leakage and privacy protection. Privacy-preserving techniques like differential privacy (Dwork, 2006) and federated learning (Kairouz et al., 2021) enable organizations to unlock collective insights across silos—think hospitals training shared cancer-detection models without ever moving patient data—turning privacy from a legal checkbox into a competitive advantage. In federated-learning scenarios, the training data are distributed across multiple clients, and any sharing or transmission between clients—or between a client and the central server, if one exists—is forbidden, thereby protecting data privacy to some extent. This protection is often further strengthened by differential privacy, with a trade-off between accuracy and privacy.
Recently, there have been a few studies considering federated learning for missing data imputation (Li et al., 2021; Abbasi et al., 2023; Lian et al., 2025; Fang et al., 2025a; Li et al., 2025c), but there are still a few challenges.
  • It is difficult to extend nonparametric imputation methods such as kNN and tree-based methods to federated learning scenarios (Zhang et al., 2024a; Li et al., 2020). For example, regarding kNN, we cannot directly compute the distance between two data points located on different clients.
@@ 10.4 Pretrained and Universal Models for Missing Data Imputation
In recent years, pretrained models, such as BERT (Devlin et al., 2019), CLIP (Radford et al., 2021), and GPTs (Achiam et al., 2023), have become the centerpiece of AI research, delivering breakthrough performance on a wide spectrum of real-world problems. A pretrained model is a large neural network whose parameters have already been learned from massive, general-domain data—text, images, audio, protein sequences, or code—so it arrives packed with broad patterns and world knowledge; instead of training from scratch, practitioners simply fine-tune or prompt it with a handful of examples, slashing data, compute, and time by orders of magnitude. It is natural to consider using pretrained models in missing data imputation or even establishing a universal model for missing data imputation. Here, a universal model for missing data imputation refers to one that, once well trained, can impute missing values for any dataset—or at least a broad class of datasets from different domains—without requiring further tuning or with only light fine-tuning.
As discussed in Section 5.7, a few scholars have explored the use of LLMs to assist in missing data imputation (Huang and Hao, 2024; Hayat and Hasan, 2024; Fang et al., 2025b) and achieved some inspiring results in general missing data imputation tasks. However, there are still limitations:
@@ 11 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No.62376236 and the General Program of Natural Science Foundation of Guangdong Province under Grant No.2024A1515011771.